[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun May 21, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon May 22 05:00:30 UTC 2017


HSU Shift Log for Mon May 22 2017 01:00:24 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: The 00z TWO had to be sent out by HFO because of AWIPS  
transmission issues locally. For the 6z it was resolved. The graphics had  
no issues because they use the ATCF copy (I believe).

Other Techincal issues: Himawari data is down. Also the Indian Ocean  
satellite data is still out-- unsure if an ETA for that data to return has  
been announced. The floor by the printers is unsteady-- seems hazardous.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170522/bcbc1985/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list